OpenStudy (anonymous):

X^2+7x+9=0 Quadratic formula to solve each equation

OpenStudy (anonymous):

The formula for the quadratic formula is \[\frac{ -2\pm \sqrt{b ^{2}-2(a)(c)} }{ 2(a) }\] \[\frac{ -7\pm \sqrt{7^{2}-2(1)(9)}}{2*1}\]

OpenStudy (anonymous):

The "-2" should be "-b"

OpenStudy (anonymous):

could you crow for me how you did that

OpenStudy (anonymous):

\[\frac{ -b \pm \sqrt{b ^{2}-2(a)(c)} }{ 2a }\] So that is your equation you will be using. You are given X^2+7x+9=0 So a=1, b=7, and c = 9 You get these numbers by ignore the X^2 and X